How to find error if it is more than 10%? Here is the question. ...-:By what percentage should the pressure of a given mass of a gas be increased,so as to decrease its volume by 10% at a constant temperature OPTIONS-; A) 11.1% B) 10.1% C) 9.1% and D) 8.1%

Grade:11

1 Answers

Vikas TU
7 years ago
At constant temp. that is at isothermal.
P1V1 = constant.
so we could say here that Pressure would be inversely proprtional to the volume from here Boyl’s law.
Let 
PV = k …..........(1)
after dcreasing pressure by 10% that is V net = V – 0.1V = 0.9 V this is remaining Volume.
Now,
P’ * 0.9V = k...............(2)
divid ethe both eqn.
P/0.9P’ =  1
or
P’ = P/0.9 =  > 10P/9 = 1.111P
% increase in Pressure = > (1.11P – P)/P => 0.11P/P = > 0.11
= 11%